OpenStudy (anonymous):

Describe the transformation of the graph of f into the graph of g as either a horizontal or vertical stretch. f(x) = sqrt(x) g(x) = sqrt(0.5x)

OpenStudy (michele_laino):

let's consider the point x=4, then we have: f(4)=2 g(4)=sqrt(2) so the ratio, is: f(4)/g(4)=sqrt(2) furthermore, if x=9 we can write: f(9)=3, and g(9)=3/(sqrt(2) and the ratio is f(9)/g(9)=sqrt(2)

OpenStudy (michele_laino):

so, we have a vertical stretch

OpenStudy (anonymous):

Ok would that be all for the question?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thank you